# Photometry and Spectroscopy of SN 2024pxl: A Luminosity Link Among Type Iax Supernovae

### v1

@@ 5.1.3 Post-maximum and late-time spectra
5.1.4 NIR spectra
Figure 16 displays the NIR spectral evolution of SN 2024pxl from 6.7 to 42 days post maximum. The first two NIR spectra of SN 2024pxl show underdeveloped Fe II features with evidence for the Mg II 10,952 line. Signatures of C I 10,693 and C I 11,754 might be identifiable in the early NIR spectra of SN 2024pxl. As the ejecta cool and expand, the spectrum-formation region recedes to lower velocities and lines become somewhat less blended, with the Co II lines increasing in prominence over time. Our spectral sequence shows that Fe II and Co II lines dominate the NIR. These lines are distinctive features in the NIR spectra of all Type Iax SNe (Stritzinger et al., 2014), and Figure 17 displays this similarity in NIR spectra from bright to faint objects.
The NIR spectrum of SN 2024pxl at 23.4 days shows more similarities with the higher luminosity SN 2005hk (Kromer et al., 2013) than with the slightly less luminous SN 2019muj (Barna et al., 2021) and much less luminous SN 2010ae (Stritzinger et al., 2014). In particular, the Mg II absorption, the features at 12,000 Å and 12,700 Å, and the prominence of the Co II lines in the 16,000 Å region are differentiating features. Where SN 2024pxl resembles SN 2005hk, SN 2019muj instead resembles SN 2010ae, mainly owing to the lower velocities that allow individual lines to be separated more clearly. This NIR spectral comparison provides additional support that intermediate-luminosity objects of the class (like SNe 2024pxl and 2019muj) are transitional between high- and low-luminosity objects.
5.2 Velocity Evolution
@@
SN 2024pxl exhibits lower velocities than other SNe shown in Figure 19 except SN 2008ha. The prominent Si II line is overlapped by a continuously growing Fe II feature in the red wing of the Si II line. As a result, we systematically underestimate the velocities associated with this line over time by considering it as a single absorption feature. In the case of SN 2020udy, Maguire et al. (2023) showed that the Si II line is contaminated by the Fe II line before maximum light, although Si II is present in post-maximum spectra. This might be why a significant drop is observed in the Si II velocities post-maximum in SN 2024pxl. It is likely that more careful modeling of these lines will be required to obtain a more accurate comparison of the Si II line-velocity evolution across Type Iax SNe.
@@
The Si II and C II line velocities of SN 2024pxl at peak are 3800 km s-1 and 3500 km s-1, respectively. The measured Fe II velocities are 7100, 7060, and 5950 km s-1 at epochs , 5.8, and 1.3 days since Bmax, respectively. At a similar epoch, the estimated velocities using Fe II lines are higher than the Si II and C II line velocities, indicating significant mixing in the burning products (Phillips et al., 2007). Kwok et al. (2025) also find evidence of mixing in SN 2024pxl from the emission lines of IGEs and intermediate-mass elements (IMEs), which are all centrally peaked and have similar velocity offsets and widths. The measured Co II line velocities in the NIR region for SN 2024pxl at an epoch of 23.4 days are 3030, 2730, and 2750 km s-1 for Co II 15,759, Co II 16,064, and Co II 16,361, respectively (consistent with one another), and slightly higher than those reported by Barna et al. (2021) at an epoch of 26.5 days past Bmax for SN 2019muj. At a comparable epoch, velocities of Co II lines in SN 2010ae (Stritzinger et al., 2014) are lower than in SN 2024pxl.
6 Summary
@@
SN 2024pxl is a nearby Type Iax SN with an early detection that enabled precise constraints on its rise time. With a peak absolute magnitude of mag, SN 2024pxl is an intermediate-luminosity Type Iax SN similar to SN 2019muj. The color evolution of SN 2024pxl aligns closely with that of other Type Iax SNe. Utilizing Swift UV coverage around peak brightness, we estimate the contribution of UV flux to the bolometric luminosity and find that SN 2024pxl’s bolometric luminosity falls between those of SNe 2002cx and 2019muj, placing it in a unique intermediate position in the luminosity range of Type Iax SNe. Analytical modeling of the pseudobolometric light curve indicates the synthesis of M⊙ of 56Ni and an ejected mass of M⊙. The explosion parameters calculated through analytical modeling of the pseudobolometric light curve are consistent with the N1def model from Fink et al. (2014), a weak deflagration of a near- WD with one ignition point. This is consistent with the results of Kwok et al. (2025) that SN 2024pxl displays spectroscopic similarity to the predictions of the N1def model.
@@
Our spectral comparisons also suggest that SN 2024pxl is a transitional object between high-luminosity and low-luminosity Type Iax SNe. It exhibits many similarities to SN 2019muj and high-luminosity Type Iax SNe. In the early spectral sequence, we also detect C II that is seen more prominently in less luminous objects. Additionally, the late-time spectra of SN 2024pxl show narrow emission lines around 7300 Å whereas the high-luminosity objects (e.g., SNe 2008ge and 2020udy) display broader features.
@@
In the NIR, SN 2024pxl shows signatures from Mg II, Fe II, and Co II and it resembles the high-luminosity Type Iax SN 2005hk, while SN 2019muj more closely resembles the low-luminosity Type Iax SN 2010ae. This is mainly due to more line blending in SN 2024pxl compared with SN 2019muj, caused by wider line widths. Despite the higher velocities indicated by the wider line widths in SN 2024pxl, the evolution of the photospheric velocity of SN 2024pxl shows systematically lower velocities relative to other comparison SNe, except for the very low-luminosity SN 2008ha. Kwok et al. (2025) find potential evidence for bulk motion on the order of a few 100 km s-1 in the ejecta of SN 2024pxl from measurements of MIR forbidden lines, which would impact the true vs. observed photospheric velocity. Given the spectroscopically and photometrically intermediate position of SN 2024pxl between SN 2019muj and brighter Type Iax SNe, we suggest that velocity measurements of the widths of the Co II lines may be better correlated with luminosity than the photospheric velocity. These spectral properties provide evidence that SN 2024pxl is a transitional Type Iax SN.
@@
7 acknowledgments
We thank Stéphane Blondin for insightful and helpful comments which improved the manuscript.
@@
A major upgrade of the Kast spectrograph on the Shane 3 m telescope at Lick Observatory, led by Brad Holden, was made possible through generous gifts from the Heising-Simons Foundation, William and Marina Kast, and the University of California Observatories. Research at Lick Observatory is partially supported by a generous gift from Google. We would like to express our gratitude to the Lick Observatory staff for their support. Shane 3 m observations were conducted on the stolen land of the Ohlone (Costanoans), Tamyen and Muwekma Ohlone tribes. We gratefully acknowledge the usage of native lands for our science. Observations from coauthor A.J.M. were made under the aegis of the ASTRAL (Astronomy/STEM Alliance with Lick Observatory) consortium, supported by a generous grant from the Gordon and Betty Moore Foundation (PI B. Macintosh).
@@ Appendix A Optical photometry and spectroscopy of SN 2017drh
We present the optical Las Cumbres Observatory photometry and FLOYDS spectroscopy for SN 2017drh, which exploded in the same host galaxy, NGC 6384, as SN 2024pxl. The spectra were reduced using the FLOYDS reduction pipeline (Valenti et al., 2014). The photometric observations were taken as part of the Global Supernova Project (GSP) collaboration. These data were reduced with lcogtsnpipe (Valenti et al., 2016b), a PyRAF-based image-reduction pipeline. In Figure 20, the LCO data are shown with the BayeSN light curve fits that also included data from Stahl et al. (2019).

### v2

@@ V.1.3 Post-maximum and late-time spectra
V.1.4 NIR spectra
Figure 16 displays the NIR spectral evolution of SN 2024pxl from 6.7 to 42.5 days post maximum. The first two NIR spectra of SN 2024pxl show underdeveloped Fe II features with evidence for the Mg II 10,952 line. Possible signatures of C I 10,693 and C I 11,754 are suggested in the early NIR spectra of SN 2024pxl. As the ejecta cool and expand, the spectrum-formation region recedes to lower velocities and lines become somewhat less blended, with the Co II lines increasing in prominence over time. Our spectral sequence shows that Fe II and Co II lines dominate the NIR. These lines are distinctive features in the NIR spectra of all Type Iax SNe (Stritzinger et al., 2014), and Figure 17 displays this similarity in NIR spectra from bright to faint objects.
The NIR spectrum of SN 2024pxl at 23.4 days shows more similarities with the high-luminosity SN 2005hk (Kromer et al., 2013) than with the intermediate-luminosity SN 2019muj (Barna et al., 2021) and low-luminosity SN 2010ae (Stritzinger et al., 2014). In particular, the Mg II absorption, the features at 12,000 Å and 12,700 Å, and the prominence of the Co II lines in the 16,000 Å region are differentiating features. Where SN 2024pxl resembles SN 2005hk, SN 2019muj instead resembles SN 2010ae, mainly owing to the lower velocities that allow individual lines to be separated more clearly. This NIR spectral comparison provides additional support that intermediate-luminosity objects of the class (like SNe 2024pxl and 2019muj) are transitional between high- and low-luminosity objects.
V.2 Velocity Evolution
@@
SN 2024pxl exhibits lower velocities than other SNe shown in Figure 19 except low-luminosity SN 2008ha. The prominent Si II line is overlapped by a continuously growing Fe II feature in the red wing of the Si II line. As a result, we systematically underestimate the velocities associated with this line over time by considering it as a single absorption feature. In the case of high-luminosity SN 2020udy, Maguire et al. (2023) showed that the Si II line is contaminated by the Fe II line before maximum light, although Si II is present in post-maximum spectra. This might be why a significant drop is observed in the Si II velocities post-maximum in SN 2024pxl. It is likely that more careful modeling of these lines will be required to obtain a more accurate comparison of the Si II line-velocity evolution across Type Iax SNe.
@@
The Si II and C II line velocities of SN 2024pxl at peak are 3800 km s-1 and 3500 km s-1, respectively. The measured Fe II velocities are 7100, 7060, and 5950 km s-1 at epochs , 5.8, and 1.3 days since Bmax, respectively. At a similar epoch, the estimated velocities using Fe II lines are higher than the Si II and C II line velocities, indicating significant mixing in the burning products (Phillips et al., 2007). Kwok et al. (2025) also find evidence of mixing in SN 2024pxl from the emission lines of IGEs and intermediate-mass elements (IMEs), which are all centrally peaked and have similar velocity offsets and widths. The measured Co II line velocities in the NIR region for SN 2024pxl at an epoch of 23.4 days are 3030, 2730, and 2750 km s-1 for Co II 15,759, Co II 16,064, and Co II 16,361, respectively (consistent with one another), and slightly higher than those reported by Barna et al. (2021) at an epoch of 26.5 days past Bmax for intermediate-luminosity SN 2019muj. At a comparable epoch, velocities of Co II lines in low-luminosity SN 2010ae (Stritzinger et al., 2014) are lower than in SN 2024pxl.
VI Summary
@@
SN 2024pxl is a nearby Type Iax SN with an early detection that enabled precise constraints on its rise time. With a peak absolute magnitude of mag, SN 2024pxl is an intermediate-luminosity Type Iax SN similar to SN 2019muj. The color evolution of SN 2024pxl aligns closely with that of other Type Iax SNe. Utilizing Swift UV coverage around peak brightness, we estimate the contribution of UV flux to the bolometric luminosity and find that SN 2024pxl’s pseudo bolometric luminosity falls between those of SNe 2002cx and 2019muj, placing it in a unique intermediate position in the luminosity range of Type Iax SNe. Analytical modeling of the integrated SED light curve indicates the synthesis of M⊙ of 56Ni and an ejected mass of M⊙. The explosion parameters calculated through analytical modeling of the integrated SED light curve are consistent with the N1def model from Fink et al. (2014), a weak deflagration of a near- WD with one ignition point. This is consistent with the results of Kwok et al. (2025) that SN 2024pxl displays spectroscopic similarity to the predictions of the N1def model.
@@
Our spectral comparisons also suggest that SN 2024pxl is a transitional object between high-luminosity and low-luminosity Type Iax SNe. It exhibits many similarities to intermediate-luminosity SN 2019muj and high-luminosity Type Iax SNe. In the early spectral sequence, we also detect C II that is seen more prominently in less luminous objects. Additionally, the late-time spectra of SN 2024pxl show narrow emission lines around 7300 Å whereas the high-luminosity objects (e.g., SNe 2008ge and 2020udy) display broader features.
@@
In the NIR, SN 2024pxl shows signatures from Mg II, Fe II, and Co II and it resembles the high-luminosity Type Iax SN 2005hk, while intermediate-luminosity Type Iax SN 2019muj more closely resembles the low-luminosity Type Iax SN 2010ae. This is mainly due to more line blending in SN 2024pxl compared with SN 2019muj, caused by wider line widths. Despite the higher velocities indicated by the wider line widths in SN 2024pxl, the evolution of the photospheric velocity of SN 2024pxl shows systematically lower velocities relative to other comparison SNe, except for the very low-luminosity SN 2008ha. Kwok et al. (2025) find potential evidence for bulk motion on the order of a few 100 km s-1 in the ejecta of SN 2024pxl from measurements of MIR forbidden lines, which would impact the true vs. observed photospheric velocity. Given the spectroscopically and photometrically intermediate position of SN 2024pxl between SN 2019muj and brighter Type Iax SNe, we suggest that velocity measurements of the widths of the Co II lines may be better correlated with luminosity than the photospheric velocity. These spectral properties provide evidence that SN 2024pxl is a transitional Type Iax SN.
@@
Figure 20 shows the distribution of Type Iax SNe in terms of their peak absolute magnitude and light-curve decline rate in the band. The data are taken from Singh et al. (2023) and references therein, augmented with the recently studied Type Iax events SN 2022eyw (Das et al., 2025), SN 2022xlp (Bánhidi et al., 2025), SN 2025qe (Magee et al., 2025), and SN 2024pxl. Singh et al. (2023) suggested that bright (high-luminosity) Type Iax SNe have mag, while faint (low-luminosity) ones have mag, and further proposed a negative linear correlation for the bright population and a positive correlation for the faint population. SN 2024pxl lies near the faint boundary of the bright Type Iax group, whereas SN 2022eyw is a luminous member located well within the bright population. In contrast, SNe 2022xlp and 2025qe exhibit intermediate luminosities, bridging the bright and faint subclasses. The newly added events follow the overall trend, with SN 2024pxl positioned close to the lower end of the bright group, underscoring its transitional nature. As additional Type Iax SNe spanning the full luminosity range become available, these correlations can be tested more robustly.
@@
VII acknowledgments
We thank the referee for the careful review and valuable suggestions, which have helped refine the manuscript. We also thank Stéphane Blondin for helpful comments which improved the manuscript.
@@
This study employs observations obtained with the Hobby-Eberly Telescope, which is a joint project of the University of Texas at Austin, the Pennsylvania State University, Ludwig-Maximilians-Universität München, and Georg-August-Universität Göttingen. The HET is named in honor of its principal benefactors, William P. Hobby and Robert E. Eberly. The Low Resolution Spectrograph 2 (LRS2) was developed and funded by the University of Texas at Austin McDonald Observatory and Department of Astronomy and by Pennsylvania State University. We thank the Leibniz-Institut für Astrophysik Potsdam (AIP) and the Institut für Astrophysik Göttingen (IAG) for their contributions to the construction of the integral field units.
KAIT and its ongoing operation were made possible by donations from Sun Microsystems, Inc., the Hewlett-Packard Company, AutoScope Corporation, Lick Observatory, the U.S. NSF, the University of California, the Sylvia & Jim Katzman Foundation, and the TABASGO Foundation. A major upgrade of the Kast spectrograph on the Shane 3 m telescope at Lick Observatory, led by Brad Holden, was made possible through generous gifts from the Heising-Simons Foundation, William and Marina Kast, and the University of California Observatories. Research at Lick Observatory is partially supported by a generous gift from Google. We would like to express our gratitude to the Lick Observatory staff for their support. Shane 3 m observations were conducted on the stolen land of the Ohlone (Costanoans), Tamyen and Muwekma Ohlone tribes. We gratefully acknowledge the usage of native lands for our science. Observations from coauthor A.J.M. were made under the aegis of the ASTRAL (Astronomy/STEM Alliance with Lick Observatory) consortium, supported by a generous grant from the Gordon and Betty Moore Foundation (PI B. Macintosh).
@@ Appendix A Optical photometry and spectroscopy of SN 2017drh
We present the optical Las Cumbres Observatory photometry and FLOYDS spectroscopy for SN 2017drh, which exploded in the same host galaxy, NGC 6384, as SN 2024pxl. The spectra were reduced using the FLOYDS reduction pipeline (Valenti et al., 2014). The photometric observations were taken as part of the Global Supernova Project (GSP) collaboration. These data were reduced with lcogtsnpipe (Valenti et al., 2016b), a PyRAF-based image-reduction pipeline. In Figure 21, the LCO data are shown with the BayeSN light curve fits that also included data from Stahl et al. (2019b).